### v1

# Curriculum learning based pre-training using Multi-Modal Contrastive Masked Autoencoders

Abstract
In this paper, we propose a new pre-training method for image understanding tasks under Curriculum Learning (CL) paradigm which leverages RGB-D. The method utilizes Multi-Modal Contrastive Masked Autoencoder and Denoising techniques. Recent approaches either use masked autoencoding (e.g., MultiMAE [6]) or contrastive learning(e.g., Pri3D [45], or combine them in a single contrastive masked autoencoder architecture such as CMAE [47] and CAV-MAE [33]. However, none of the single contrastive masked autoencoder is applicable to RGB-D datasets. To improve the performance and efficacy of such methods, we propose a new pre-training strategy based on CL. Specifically, in the first stage, we pre-train the model using contrastive learning to learn cross-modal representations. In the second stage, we initialize the modality-specific encoders using the weights from the first stage and then pre-train the model using masked autoencoding and denoising/noise prediction used in diffusion models. Masked autoencoding focuses on reconstructing the missing patches in the input modality using local spatial correlations, while denoising learns high frequency components of the input data. Our approach is scalable, robust and suitable for pre-training with limited RGB-D datasets. Extensive experiments on multiple datasets such as ScanNet, NYUv2 and SUN RGB-D show the efficacy and superior performance of our approach. Specifically, we show an improvement of +1.0% mIoU against Mask3D on ScanNet semantic segmentation. We further demonstrate the effectiveness of our approach in low-data regime by evaluating it for semantic segmentation task against the state-of-the-art methods.
1 Introduction
Self-supervised learning (SSL) has paved the way for improving performance in various computer vision tasks such as object recognition [38, 87, 9], semantic segmentation [43, 29, 53] and depth estimation [35] by leveraging extensive unlabeled data and large models such as Vision Transformers (ViT) [26, 57]. There are two self-supervised frameworks commonly used to pre-train large vision models. First, contrastive learning [17, 89, 19, 34] pre-trains a model by maximizing the similarity between different augmentations of an image while minimizing the similarity to the other data samples. It essentially learns discriminative representations by enforcing invariance to data augmentations. Second family of SSL is Masked Image Modeling (MIM) which is inspired by the recent success of masking in the NLP domain [23]. The seminal work in MIM is Masked Autoencoders (MAE) [38] which masks out high number of patches in the input before passing them into a Transformer encoder. Then, a lightweight decoder is used to reconstruct the masked patches in the input. Such reconstruction task provides the model with understanding of the data beyond low-level image statistics. It has achieved state-of-art performance in a variety of image recognition tasks.
Recently, Masked autoencoders have been extended to pre-train transformers using multi-modal data for tasks such as video-text retrieval [56, 83], image-text retrieval [68] and video-audio-text understanding [1]. However, utilizing MAEs for dense prediction tasks like semantic segmentation and depth estimation requires understanding beyond image level statistics. With the availability of RGB-D or Time-of-Flight (ToF) cameras, various approaches [6, 43, 31, 27] have been developed to improve the dense prediction tasks using multi-modal datasets [4, 15, 22, 63, 77]. A few methods [6, 43] also leverage RGB-D datasets to pre-train ViTs backbones under MAE paradigm. One such work is MultiMAE [6] which pre-trains the vision transformers using RGB, depth and segmentation masks in a multi-task MAE fashion. However, MultiMAE requires semantic segmentation labels during pre-training and access to multiple modalities during fine-tuning for downstream tasks. This method learns cross-modal relationships using a cross-attention module but under the paradigm of MAE to help reconstruct the masked patches in each modality. Another work in this line is Mask3D [43] which embeds 3D priors in ViT backbones to learn structural and geometric priors by reconstructing the masked patches in the depth input using RGB-D data. However, it doesn’t learn cross-modal relationships to capture more local context beyond 3D priors. Some prior works [20, 86, 92, 44] have also used contrastive learning for high-level scene understanding and low-level point matching [8, 91].
Motivated by this, we believe contrastive learning and Masked autoencoders can be complementary to each other as both can provide unique data representations by extracting different discriminative features from RGB-D datasets. However, MAEs alone, only focus on local spatial statistics for reconstruction. How about learning high frequency components of the data? This further motivates us to explore another learning mechanism. Recently, we have seen a remarkable success in image and video generation [41, 40, 42] using diffusion models and this success comes from the denoising technique. Inspired by this, we introduce a noise prediction loss with the masked autoencoder reconstruction objective. To do so, we add a Gaussian noise in the input and pre-train the model to predict the masked patches as well as the noise. We hypothesize that the denoising would encourage the model to extract high-frequency features from the input.
The main challenge we’re addressing in this work is how to effectively combine such methods in one single pre-training framework especially for RGB-D datasets. We have seen recent works that combine instance-level contrastive learning and masked autoencoders and leverage multi-modal data in one single framework for image recognition [47], video action recognition [60], video-text [61] and audio-video [33]. However, such methods are not suitable for RGB-D datasets.
Dense prediction tasks such as semantic segmentation and depth estimation require understanding beyond image-level representation. We hypothesize that the instance-level contrastive learning would not benefit the aforementioned downstream tasks when leveraging RGB-D datasets for pre-training. So, to capture useful discriminative local context features using these modalities, patch-level or pixel-level contrastive learning would be more beneficial. However, it is not easy to combine patch-level or pixel-level contrastive learning with Masked Autoencoders in a single framework because of the masking step.
Based on the above challenges, we propose a new strategy based on these self-supervised frameworks to pre-train ViTs using RGB-D datasets. Our curriculum consists of two stages, 1) contrastive learning at pixel or patch-level to align RGB and depth modalities to capture local context for cross-modal relationships 2) MIM + Denoising would encourage the encoder to learn discriminative low-frequency and high-frequency features respectively. Indeed, one can change the ordering in the curriculum, but we empirically found that learning an easier task i.e., contrastive learning is much more beneficial and improves the downstream performance. Furthermore, from the figure 1, contrastive learning converges faster compared to MIM + Denoising, making it a relatively easier task in comparison. It would also facilitate the learning to next round of pre-training (MIM + Denoising) in the curriculum. Lastly, we pre-train ViT-B on ScanNet [22] and SUN RGB-D [77] with our curriculum learning based pre-training and evaluate it on multiple downstream tasks.
Figure 1: Convergence of each learning phase of our curriculum on ScanNet RGB-D dataset. (Left) InfoNCE loss converges faster than (Right) Masked Image Modeling and Denoising making it a relatively easier task.
2 Related Work
Curriculum learning.
Humans learn information effectively when there’s a structured order in place, not by randomly presenting data samples to them. For example, there is a particular curricula for students to follow in the school. Inspired by this, the term curriculum learning in ML was first proposed in [10], providing easy-to-hard training strategies for a model. This allows the model to converge faster and achieve better performance. There has been a number of works proposing different curriculum learning strategies for various computer vision tasks such as object localization [48, 75, 79], object detection [18, 74, 82, 55] and other applications [80, 72, 36, 37, 32, 67, 51].
Contrastive learning based pre-training.
SSL consists of designing a pre-text task to learn discriminative representations from unlabeled data. Earlier work on constructing pre-text tasks include colorization [90], jigsaw puzzle [65], and predicting the rotation angle of image [30]. Currently, the two popular approaches for self-supervised learning are constrastive learning and masked image modeling. Contrastive learning groups similar data samples in the feature space while pushing away the dissimilar ones. SimCLR [17] uses contrastive learning to maximize the similarity between different augmentations of an image. MoCo [89] uses dictionary as a queue to facilitate contrastive learning. BYOL [34] uses two networks termed as online and target, to learn from each other by using augmented views of an image. On the other-hand, there are clustering based unsupervised methods that either use k-means assignments as pseudo-labels [11] or enforce similarity between the cluster assignments of different augmented views of the image [12].
Masked Autoencoder based pre-training.
Masked Image Modeling (MIM) has received a lot of attention for pre-training vision models since the advent of Vision Transformers (ViT). The goal of MIM is to learn discriminative features for downstream tasks by reconstructing the missing patches in the input. Masked autoencoding for vision tasks is inspired by the success of methods such as BERT [23] and GPT [69] in NLP domain. Several MIM approaches have been proposed with various reconstruction objectives based on pixels [5, 26, 16, 28, 38, 87, 50], dVAE tokens [9, 93] and features such as histogram of oriented gradients (HOG) or deep neural network embeddings [84, 7]. Specifically, MAE [38] uses an asymmetric encoder-decoder architecture based on ViT to pre-train models using masked autoencoding. It first masks large portions of an image and pass the unmasked patches into encoder followed by a decoder for reconstructing the missing patches. The masking accelerates the pre-training of ViT. Contrastive learning is different from MAE in the sense that MAE learns local spatial dependencies, while contrastive approaches learn representations such as invariance to different augmentations. This steers us to hypothesize that these two SSL frameworks learn complimentary features and therefore we combine them in a curriculum manner to get the best out of each. Indeed, one can combine them in a single framework [47, 33], but our empirical experiments show poor performance of such methods on RGB-D datasets.
Multi-Modal Learning.
It involves learning from multiple modalities such as images and text [2, 52, 88, 14, 46, 21, 78], video and audio [3, 49, 62, 66], video, text and audio [1], and RGB and depth images [31]. Learning can be performed either using modality-specific encoders or shared ones. In terms of MIM, MultiMAE [6] pre-trains the ViT using RGB, depth and semantic segmentation while reconstructing the missing patches in these modalities. It also uses additional modality during the fine-tuning. Mask3D [43] uses RGB and depth modalities under MIM paradigm to encode the 3D priors by reconstructing the missing patches in the depth input.
Denoising.
It involves reconstructing clean data given a noisy input. Denoising has recently revolutionized generative models that produce high quality images using diffusion modeling [24, 41]. It consists of forward Gaussian diffusion process and backward generation process. In the forward process, it adds Gaussian noise to the sample while in reverse process, it learns to generate high quality image from the Gaussian noise. This has been used for generating text-conditioned images [64, 70, 73] and videos [76, 40, 42]. The success of denoising hasn’t been utilized in representation learning and in this work, we leverage the denoising process to investigate the effectiveness of representations learned from RGB-D data.
3 Method
Our approach is a simple combination of constrastive learning, masked autoencoding and denoising used in diffusion models, designed in a curriculum for pre-training vision transformers using RGB-D data. Our goal is to learn representations from different learning mechanisms that can complement each other. The contrastive learning enhances the cross-modal representations via pixel-to-pixel level or patch-level learning, while Masked autoencoding understands the local spatial correlations, while providing 3D priors as a guidance for the RGB input. Moreover, the noise prediction focuses on the high-frequency information of the input data. We first give an overview of our approach and then explain the learning objectives i.e, contrastive, masked image modeling and noise prediction. Finally, we describe the overall pre-training curriculum pipeline.
3.1 Overview
Given a batch of $n$ RGB-D pair $\{\mathbf{x}^{rgb},\mathbf{x}^{depth}\}_{i=1}^{n}$, we pre-train ViT [26] using the combination of self-supervised frameworks in a curriculum learning manner. Our curriculum learning based pre-training consists of two stages. In the first stage, we pre-train our model using an easier task which is to learn cross-modal representations to capture correspondence using constrastive learning either at pixel-to-pixel level or patch level. In the second stage, we first initialize the encoders using the weights learned in the previous stage. Then, we use multi-modal masked autoencoder to learn more fine-grained features beyond low-level image statistics to reconstruct the missing patches in the depth input. To pre-train using masked autoencoding, we first split the RGB-D input in $T=(h/p)\times(w/p)$ patches of size $p\times p$: $\mathbf{x}_{i,\text{patch}}^{rgb},\mathbf{x}_{i,\text{patch}}^{depth}\in\mathbb{R}^{T\times p\times p\times 3}$ to pass them to the encoders. We use modality-specific encoders instead of joint encoder in the second stage. We then randomly mask patches in both inputs by keeping a percentage of patches and obtain two masks $\mathbf{M}_{i}^{rgb},\mathbf{M}_{i}^{depth}$. Following MAE [38], we pass the unmasked patches to the encoders to obtain the latent representations. The representations are then concatenated with the learnable masked tokens which are essentially the placeholders for these tokens. A lightweight decoder is used to reconstruct the masked patches. Furthermore, we also add denoising to learn the high frequency information of data. Figure 2 shows the overview of our approach.
Figure 2: Overview of our curriculum learning based pre-training. In the first stage, we pre-train the ViT-B encoder using Contrastive Learning to align the RGB and the depth patches. In the second stage, we initialize the modality-specific ViT-B encoders with the stage-1 weights and pre-train them using Multi-modal Masked autoencoding and Denoising to reconstruct the masked patches in depth input and the noise in the unmasked patches respectively.
3.2 Contrastive learning
To learn the correspondence between RGB and depth modalities, we use contrastive learning to discriminate the instances at patch-level or pixel-level. There has been a few works [58, 54] using constrastive learning to learn correspondence between RGB and pointclouds. The depth values can be indirectly related to the pointclouds and projected to 3D pointclouds using camera parameters. Therefore, we can also leverage the same framework of pixel-to-point [58] to learn the correspondence between RGB and depth. We experimented with both pixel-level and patch-level contrastive learning and found no empirical differences in the downstream performance. Specifically, for a given RGB $\mathbf{x}_{i}^{rgb}$ and depth $\mathbf{x}_{i}^{depth}$, we pass them through the shared ViT encoder to get the latent representations $\mathbf{z}^{rgb}_{i},\mathbf{z}_{i}^{depth}$. We use Conv2d as patch projection layer. Following [81], we use InfoNCE loss to pre-train the encoder with a fixed temperature $\tau$:
$\mathcal{L}_{\mathrm{PNCE}}=-\frac{1}{N}\sum_{i=1}^{N}{\rm log}\left[\frac{{\rm exp}(s_{i,i}/\tau)}{\sum_{k\neq i}{\rm exp}(s_{i,k}/\tau)+{\rm exp}(s_{i,i}/\tau)}\right]$ (1)
where $s_{i,j}=\|\mathbf{z}^{rgb}_{i}\|^{T}\|\mathbf{z}_{i}^{depth}\|$. $\|\mathbf{z}^{rgb}_{i}\|$ and $\|\mathbf{z}_{i}^{depth}\|$ correspond to the encoder features for RGB and depth respectively for a patch or pixel $i$. In the experiments, we apply the InfoNCE loss between $\mathbf{z}^{rgb}_{i}$ , $\mathbf{z}_{i}^{depth}$ , $\mathbf{z}^{depth}_{i}$ and $\mathbf{z}_{i}^{rgb}$, and take the average of the losses. This loss aligns the RGB and depth patch by maximizing the similarity using the latent representations, while minimizing the similarity with other depth patches for a given RGB patch. Indeed, an instance-level contrastive learning can be used but it would fail to capture local discriminative features across modalities as the task becomes much easier by only learning high-level semantics.
3.3 Masked Autoencoding
The goal of masked autoencoding is to understand local spatial statistics of the input modality. To further improve the capability of our approach, we propose to pre-train the model using MAE [38] as a second learning step in our curriculum. We initialize our modality-specific encoders with stage-1 weights. Unmasked patches of two modalities are passed to the encoders to get latent embeddings. Embeddings are then concatenated with learnable masked tokens which are then passed to the lightweight decoder. We then apply depth prediction MLP head to reconstruct the missing patches. We also experiment with RGB reconstruction task but we found that empirically it doesn’t improve the downstream performance (c.f. Table 8). Following [38, 43], we normalize the target patches and compute the loss only on masked patches. Specifically, we use MSE loss which is given as:
$\mathcal{L}_{\mathrm{depth}}=\frac{1}{n}\sum_{i=1}^{n}\|\mathbf{M}_{i}^{depth}\circ(\mathbf{x}^{depth}_{i}-\hat{\mathbf{x}}^{depth}_{i})\|_{2}^{2}\vspace{-5pt}$ (2)
Figure 3: Denoising: We first add Gaussian noise to the unmasked patches. We then pass the noise level $\sigma$ through MLP, and add it to encoded tokens before passing them through the decoder for reconstruction.
3.4 Denoising
Denoising approaches has been shown to be a catalyst for models generating high quality images and videos [41, 40]. Inspired by them, we investigate whether denoising can also be used in self-supervised learning. During pre-training, we add an independent isotropic Gaussian noise to the depth modality $\mathbf{x}_{i}^{depth}\leftarrow\mathbf{x}_{i}^{depth}+\sigma_{i}^{depth}\mathbf{e}^{depth}_{i}$ with $\mathbf{e}^{depth}_{i}\sim\mathcal{N}(\mathbf{0},I)$ and $\sigma_{i}^{depth}$ uniformly sampled from an interval $[0,\sigma_{\text{max}}]$. This noisy input is masked and then passed to the depth encoder as mentioned in section 3.3 which follows MAE [38]. Before passing the encoded representations to the decoder, we also add the information about noise level $\sigma_{i}^{depth}$ to help separate the noise from the clean input, which is inspired from the denoising diffusion model [41]. To differentiate, we treat $\sigma_{i}^{depth}$ as the positional encoding to the decoder. We first produce a 1D sinusoidal embedding of $\sigma_{i}^{depth}$ which is then passed to the MLP with ReLU activation to produce a learnable embedding $\mathbf{p}_{i}^{depth}\in\mathbb{R}^{d}$. The dimension of this embedding matches the dimension of the representation from the encoder. We then add this embedding to the encoded embeddings as $(\mathbf{z}_{i}^{depth})\leftarrow(\mathbf{z}_{i}^{depth})+\mathbf{p}_{i}^{depth}$ before concatenating it with the RGB encoder representations and learnable masked tokens. The concatenated result is passed to the decoder to produce $\hat{\mathbf{x}}^{depth}_{i}$. We use the MSE loss and compute the denoising loss on the unmasked patches as:
$\mathcal{L}_{\mathrm{denoise}}=\frac{1}{n}\sum_{i=1}^{n}\|(1-\mathbf{M}_{i}^{depth})\circ(\sigma_{i}^{depth}\mathbf{e}^{depth}_{i}-\hat{\mathbf{x}}^{depth}_{i})\|_{2}^{2}$ (3)
Denoising would encourage the encoder to extract high frequency components which can be complementary to the MAE pre-training in the second stage of our curriculum. The computation of the denoising doesn’t bring any overhead to the pre-training compared to the vanilla MAE, since we use a lightweight MLP and compute the loss on the reconstructed unmasked patches. The decoder also reconstructs the unmasked patches which may go wasted in the vanilla MAE pre-training. We also provide an ablation study (c.f. Table 6) on the effect of each component in the denoising stage. Lastly, the figure 3 shows the overall setup of the denoising.
3.5 Pre-training Curriculum
We are now ready to present our full curriculum for pre-training ViT using RGB-D datasets. In the first stage, we pre-train the model using contrastive loss as mentioned in section 3.2.
$\mathcal{L}_{\mathrm{stage1}}=\mathcal{L}_{\mathrm{PNCE}}$ (4)
In the second stage, we pre-train the model using Masked autoencoding and denoising. Therefore, the overall stage-2 objective can be defined as:
$\mathcal{L}_{\mathrm{stage2}}=\alpha\mathcal{L}_{\mathrm{depth}}+\beta\mathcal{L}_{\mathrm{denoise}}$ (5)
Where $\alpha$, $\beta$ are the weights for each loss function during pre-training. Please see appendix on the details of these hyper-parameters.
4 Results
In this section, we show the efficacy of our approach for ViT-B backbone on semantic segmentation, instance segmentation and depth estimation. We pre-train ViT-B on the ScanNet [22] dataset which contains 2.5M RGB-D frames from 1513 training video sequences. We regularly sample every 25th frame without any filtering during pre-training following Mask3D [43]. To study the effectiveness of our pre-training approach for small scale datasets, we pre-train the model using SUN RGB-D dataset [77]. It contains 10,335 RGB-D images captured from 4 different sensors. We use the official split of train/test for pre-training and evaluation. Furthermore, we initialize our encoders using ImageNet self-supervised weights [38]. Please refer to the appendix for more details on implementations and hyper-parameters.
Methods Reconstruction task Backbone Pre-train Fine-tune Modality mIoU
Scratch - ViT-B None RGB 32.6
Pri3D [45] ViT-B ImageNet+ScanNet RGB 59.3
Pri3D [45] - ResNet-50 ImageNet+ScanNet RGB 60.2
DINO [13] - ViT-B ImageNet+ScanNet RGB 58.1
MAE [38] RGB ViT-B ImageNet RGB 64.8
MAE [38] RGB ViT-B ImageNet+ScanNet RGB 64.5
MultiMAE* [6] RGB + Depth ViT-B ImageNet+ScanNet RGB 65.1
Mask3D [43] Depth ViT-B ImageNet+ScanNet RGB 66.0
Mask3D [43] RGB + Depth ViT-B ImageNet+ScanNet RGB 65.5
Ours Depth ViT-B ImageNet+ScanNet RGB 67.0
MultiMAE [6] RGB + Depth + Segmentation ViT-B ImageNet RGB 66.4
Table 1: ScanNet Semantic Segmentation. Our approach outperforms Mask3D and other SOTA approaches that leverage RGB-D data during pre-training. * means MultiMAE was pre-trained using ScanNet by initializing the ViT-B encoder with ImageNet SSL weights.
4.1 Semantic Segmentation
We first present the results for semantic segmentation. Following MultiMAE [6], we use ConvNext architecture [59] as a segmentation decoder for our approach and the baselines. For evaluation, we use mean Intersection over Union (mIoU). Table 1 shows the performance of our approach and other baselines on ScanNet 2D semantic segmentation task. For fine-tuning, we follow Mask3D [43] and sample every 100th frame, resulting in 20,000 training images and 5000 validation images. Our approach achieves 67.0 mIoU compared to 66.0 mIoU for Mask3D and 65.1 mIoU for MultiMAE. For fair comparison, we pre-train MultiMAE with RGB and depth only. Moreover, compared to MAE pre-trained on ScanNet RGB modality, our approach improves the performance by +2.5 mIoU. Furthermore, compared to Pri3D, a multi-view contrastive learning approach, we outperforms it by +6.8 mIoU. Table 2 shows the results on SUN RGB-D 2D semantic segmentation. We observe that our method outperforms other state-of-the-art approaches. More notably, we achieve 48.1 mIoU compared to 47.4 mIoU for Mask3D and 47.1 mIoU for MultiMAE. These results demonstrate that our pre-training approach can be used for relatively small scale RGB-D datasets. Please follow supplementary materials for more results on NYUv2 semantic segmentation.
Methods Reconstruction task Backbone Pre-train Fine-tune Modality mIoU
MAE [38] RGB ViT-B ImageNet RGB 47.0
MAE [38] RGB ViT-B ImageNet+SUN RGB-D RGB 47.3
Mask3D [43] Depth ViT-B ImageNet+SUN RGB-D RGB 47.4
MultiMAE [6] RGB + Depth ViT-B ImageNet+SUN RGB-D RGB 47.1
Ours Depth ViT-B ImageNet+SUN RGB-D RGB 48.1
Table 2: Comparison with the state-of-the-art approaches on SUN RGB-D Semantic Segmentation.
4.2 Depth Estimation
In this section, we study the efficacy of our pre-trained model for depth estimation task. We use NYUv2 [63] for this downstream task and report $\delta_{1}$ on the test set which shows the percentage of pixels that have an error ratio ($\max\{\frac{\hat{y}_{p}}{y_{p}},\frac{y_{p}}{\hat{y}_{p}}\}$) below 1.25 [25]. We use DPT [71] as dense prediction head following MultiMAE for all the approaches and use the model pre-trained on ScanNet dataset for fine-tuning. The results on the depth prediction task are reported in Table 3. We can draw the same observation from the table that our approach outperforms state-of-the-art approaches. Specifically, it achieves 86.8 $\delta_{1}$ compared to 85.4 $\delta_{1}$ for Mask3D and 85.3 $\delta_{1}$ for MultiMAE. We also compare our approach with CroCo [85], a cross-view completion pre-training approach specifically designed for 3D vision tasks. In particular, we achieve an improvement of +1.2 $\delta_{1}$ over CroCo. These results again verify the generalizibility of our curriculum based pre-training strategy by transferring the representations to various downstream tasks and datasets.
Methods Reconstruction task Backbone Pre-train Fine-tune Modality $\delta_{1}$
MAE [38] RGB ViT-B ImageNet RGB 85.1
Mask3D [43] Depth ViT-B ImageNet+ScanNet RGB 85.4
CroCo [85] RGB + Depth ViT-B Habitat RGB 85.6
MultiMAE* [6] RGB + Depth + Segmentation ViT-B ImageNet RGB 83.0
MultiMAE [6] RGB + Depth ViT-B ImageNet+ScanNet RGB 85.3
Ours Depth ViT-B ImageNet+ScanNet RGB 86.8
MultiMAE [6] RGB + Depth + Segmentation ViT-B ImageNet RGB 86.4
Table 3: NYUv2 Depth Estimation. Ours approach outperforms Mask3D and MultiMAE which shows that learned representations can be transferred to the dense prediction tasks even for different dataset.* means the number are reported from CroCo [85] paper.
4.3 Instance Segmentation
We also demonstrate the effectiveness of our approach on 2D instance segmentation. We use ScanNet [22] for comparison against other baselines and use MaskRCNN [39] as the instance segmentation head on top of the ViT-B encoder. Table 4 shows the results on 2D instance segmentation. We see the same observation that our approach transfers the learned representations better by outperforming the state-of-the-art approaches including Mask3D and MultiMAE.
Methods Pre-train AP
Scratch None 12.2
ImageNet baseline Supervised ImageNet 17.6
Pri3D [45] ImageNet + ScanNet 18.3
MoCov2 [89] ImageNet + ScanNet 18.3
MAE [38] ImageNet + ScanNet 20.7
MultiMAE [6] ImageNet + ScanNet 22.4
Mask3D [43] ImageNet + ScanNet 22.8
Ours ImageNet + ScanNet 23.3
Table 4: Comparison with other state-of-the-art approaches using average precision (AP) on ScanNet 2D Instance Segmentation.
Curriculum order mIoU
Contrastive $\rightarrow$ MIM + denoising 67.0
MIM + denoising $\rightarrow$ Contrastive 64.9
Contrastive + MIM + denoising 64.2
Table 5: Effect of ordering of pre-training strategies on ScanNet 2D semantic segmentation.
4.4 Ablation Studies
Effect of curriculum order.
We study the impact of changing the order of learning in our curriculum. Table 5 shows the performance on ScanNet 2D semantic segmentation. It can be seen that our proposed curriculum strategy outperforms the other ones. More specifically, when we pre-train the model in a single framework, it performs the worst overall. We hypothesize that masking the inputs degrades the optimization of contrastive learning as it is important to learn pixel-to-pixel level or patch-level cross-modal relationships. Moreover, by pre-training using the masked image modeling first and then contrastive learning, we achieve 64.9 mIoU compared to 67.0 mIoU. This shows the importance of learning the easier task first.
Ablation study: Denoising
We consider three settings in this ablation study: 1) Without any noise 2) Input with noise only 3) full method with all the denoising components. Table 6 shows the performance of the different components of denoising loss on ScanNet 2D segmentation. It can be seen from the table that by simply adding noise in the input, the performance improves 0.5 mIoU. It is further improved by 0.8 mIoU with full denoising method that consists of noise prediction and positional encoding in the decoder.
Ablation study: Loss functions.
In Table 8, we highlight the importance of using different self-supervised frameworks in a curriculum manner. If we drop contrastive learning, our approach would only consists of Masked Image modeling and denoising, leading to a performance drop to 66.3 mIoU highlighting the importance of having constrastive learning in the first stage. Nonetheless, it is still marginally better than Mask3D (66.0 mIoU) implying that the learned representations using MIM and noise prediction are more discriminative and better transferable to the downstream task.
Method Backbone mIoU
w/out noise ViT-B 66.2
noise only ViT-B 66.7
Full ViT-B 67.0
Table 6: Ablation study on the components of the denoising method. We report the performance on ScanNet 2D semantic segmentation.
RGB ratio Depth ratio mIoU
20.0% 20.0% 65.2
20.0% 50.0% 64.9
20.0% 80.0% 65.0
50.0% 20.0% 65.5
50.0% 50.0% 65.1
80.0% 20.0% 65.7
80.0% 80.0% 67.0
Table 7: Effect of different masking ratios for RGB and depth modality on ScanNet 2D semantic segmentation.
Effect of masking ratio during pre-training.
Table 7 shows the performance on ScanNet 2D semantic segmentation by varying the masking ratio of RGB and depth modalities. It can be seen that our approach achieves the best performance with higher masking ratio for both of the modalities, which accelerates the pre-training.
Loss Backbone mIoU
MIM + Denoising ViT-B 66.3
Depth + RGB Reconstruction ViT-B 66.4
Full ViT-B 67.0
Table 8: Ablation study of different loss functions on ScanNet 2D semantic segmentation.
Effect of RGB + Depth Reconstruction.
We also study the effect of reconstructing the RGB input besides the depth modality. It can be seen from the table 8 that it doesn’t improve the performance. One plausible reason for this is that reconstructing the RGB modality is a simpler task during pre-training, given the MAE-based initialization and 3D priors are also encoded during the depth prediction.
Results on different ViT variants.
In the main manuscript, we showed the comparison of different approaches using ViT-B architecture. In Table 9, we show the performance on ScanNet 2D segmentation using ViT-L. We can infer the same observation that our approach outperforms the baselines that include MAE and MultiMAE.
Method Backbone mIoU
MAE [38] ViT-B 64.8
MultiMAE [6] ViT-B 65.1
Ours ViT-B 67.0
MAE [38] ViT-L 68.2
MultiMAE [6] ViT-L 69.3
Ours ViT-L 70.6
Table 9: Performance on ScanNet 2D semantic segmentation using variants of ViT.
Performance without ImageNet pre-training.
In Table 10, we show the performance on ScanNet 2D segmentation without initializing the models with ImageNet weights before pre-training. For this setting, we use ConvNext based decoder across all the approaches. We see a significant drop in the performance without ImageNet based initialization. Nevertheless, our approach still outperform Mask3D by +0.8 mIoU.
Method Backbone mIoU
Train from Scratch ViT-B 32.6
Mask3D [43] ViT-B 41.3
Ours ViT-B 42.1
Table 10: Without ImageNet initialization, our approach still outperforms Mask3D.
Figure 4: To show data-efficiency feature of our approach, we compare with recent state-of-the-art pre-training approaches on ScanNet 2D semantic segmentation under limited labeled data scenarios.
4.5 Low-data Regime Scenarios
We show the data-efficient nature of our pre-training strategy by fine-tuning it in limited labeled data scenarios. Figure 4 shows the performance on ScanNet 2D semantic segmentation. Our approach consistently outperform Mask3D, MAE and Pri3D across different ranges of limited labeled data. More notably, it almost achieves the same performance as MAE with only 60% of the original training set.
5 Conclusion
We propose a new curriculum learning based pre-training strategy that effectively combines powerful self-supervised learning methods in a single framework to pre-train ViTs using RGB-D datasets. Our curriculum consists of two phases of learning: (1) contrastive learning which aligns the RGB and depth patches to extract discriminative local context features, (2) Masked Image Modeling which reconstructs the masked patches in the depth modality using RGB-D input to learn spatial statistical correlations. To learn high frequency features, we introduce a denoising loss inspired by the recent success of diffusion models to predict the noise added in the input. We demonstrate the effectiveness of our approach on various downstream tasks such as semantic segmentation, depth prediction and instance segmentation. Finally, we show scalability and data-efficiency by pre-training the models using small scale RGB-D datasets.